Maria deposits
$800
into an account that pays simple interest at a rate of
5%
per year. How much interest will she be paid in the first
2
years?

Answer by Theo(13342)

simple interest is based on the original amount invested.

the formula is i = p * r * n, ifmy concept of simple interest is the same as your instructor's.

p = principal = 800
r = interest rate per time period =.05 per year.
n = number of time periods = 2 years.

i = p * r * n becomes i = 800 * .05 * 2 which is equal to 80.

with simple interest, you are earning 5% per year on your original 800.

that's 40 per year.

the formula doesn't assume anything about what you do with that 40 dollars per year that you are earning.

if you take that 40 and re-invest it into the same account, then you have a compound interest account.

the formula then becomes f = p * (1 + r) ^ n.

p is the principal
f is the future value
r is the interest rate per time period.
n is the number of time periods.

in that case, you are earning interest on interest.

f = p * (1 + r) ^ n becomes f = 800 * (1 + .05) ^ 2 which results in f = 882.

i = f - p results in interest of 82.

that extra 2 dollars is what you earned on the interest that you re-invested in the account.

your original investment is 800 for 2 years = 80 dollars.
your next investment is 40 dollars for 1 year = 2 dollars.
you earned 40 + 40 + 2 = 82 dollars in interest.

simple interest formula is i = f * r * n

compound interest formula is i = p * (1 + r) ^ n - p